OpenStudy (16shuston):

Will medal and fan. If x varies directly as y, and x=42 when y=7, find x when y=3

OpenStudy (16shuston):

all i need is the set up

Nnesha (nnesha):

equation for direct variation \[\huge\rm y=kx\] where k is constant of variation.

Nnesha (nnesha):

substitute x and y for their values solve for k

OpenStudy (anonymous):

Ok so you are going to do this: 42/7 = x/3 Now cross multiply 42*3 = 7x Solve for x

OpenStudy (16shuston):

okay that makes sense. thank you

OpenStudy (anonymous):

126 = 7x x = 18 Medal? Please